aboyles:

Two boys who live 14 miles apart start at noon to walk toward each other at rates of 3 mph and 4 mph respectively. In how many hours will they meet? 1 hr 2 hr 3 hr

Alexis1415:

Okay so basically let's focus on their walking. After one hour, one guy has walked 3mi and the other 4mi, which is in total 7 After 2 hours, one guy has walked 6mi and the other 8mi, which is 14mi So if we do it algebraically this is how we would do it: 3x +4x = 14 *x being the hour 7x = 14 x =2